OpenStudy (anonymous):

Find the equation for the ellipse with foci (1,1) and (-1,-1) and major axis of length 4 **seems to shifted an angle in the xy coordinates**

OpenStudy (anonymous):

The question say to identify the type of conic section whose equation is given and find the vertices and foci, i am unsure wheter i should mention that its seems to on an angle or do i just go ahead and write down the equation for this ellipse

Directrix (directrix):

Write the problem here exactly as the instructions appear in your book.

OpenStudy (anonymous):

identify the type of conic section whose equation is given and find the vertices and foci: Find the equation for the ellipse with foci (1,1) and (-1,-1) and major axis of length 4

OpenStudy (anonymous):

Eccentricity shall be given

OpenStudy (anonymous):

the only thing given was:Find the equation for the ellipse with foci (1,1) and (-1,-1) and major axis of length 4

OpenStudy (anonymous):

i have no problem finiding the equation, but its just that it looks tilted on the xy coordinate system, and i am wondering if i should find the angle its tilted on and how would i do so

OpenStudy (anonymous):

Easy

OpenStudy (anonymous):

\[\frac{(x-y)^2}{16}+\frac{(x+y)^2}{8}=1\]

OpenStudy (anonymous):

okay, but the titlted part, does that have any impact

OpenStudy (anonymous):

Yep it changes orientation n xy plane

OpenStudy (anonymous):

and how do i find that angle

OpenStudy (anonymous):

Angle of major axis with with horizontal is 45deg same as angle between x-y and x axis

OpenStudy (anonymous):

how did you determine that?

OpenStudy (anonymous):

how did i determine what?

OpenStudy (anonymous):

that the angle is 45 degrees

OpenStudy (anonymous):

slope of x-y=0

Directrix (directrix):

Click the link to see the ellipse with equation: (x-y)^2/(16) + (x+y)^2/(8) =1 http://www.wolframalpha.com/input/?i=%28x-y%29%5E2%2F%2816%29+%2B+%28x%2By%29%5E2%2F%288%29+%3D1

OpenStudy (anonymous):

i see its shifted at what looks like a 45 degree angle, but i am still unclear how to determine its a 45 degree angle

Directrix (directrix):

I would not worry about the angle. The instructions did not ask for it.

OpenStudy (anonymous):

great answer

OpenStudy (anonymous):

Thanx

OpenStudy (anonymous):

what about this problem:

OpenStudy (anonymous):